OpenStudy (anonymous):

1. The diagonal of a square is 8√2. Find the length of a side of the square. 2. 2x√3/2x times √8x-cubed/6

OpenStudy (anonymous):

2. \[2x \sqrt{\left(\begin{matrix}3 \\ 2x\end{matrix}\right)} \times \sqrt{\left(\begin{matrix}8x ^{3} \\6\end{matrix}\right)}\]

OpenStudy (anonymous):

Okay, I got 1. Now i just need 2.

OpenStudy (diyadiya):

Okay \[2x* \frac{\sqrt{3}}{\sqrt{2x}} * \frac{\sqrt{8x^2}}{\sqrt{6}} =\sqrt{2x}*\sqrt{3}* \frac{\sqrt{8x^2}}{\sqrt{6}} \] Since \[2x=\sqrt{2}*\sqrt{2}*\sqrt{x}*\sqrt{x}=\sqrt{2x}*\sqrt{2x}\] And\[\sqrt{6}=\sqrt{2}*\sqrt{3}\] \[\sqrt{8}x*\sqrt{x}=x \sqrt{8x}\]

OpenStudy (anonymous):

Wow.....thank you

OpenStudy (diyadiya):

Anytime :)